VISCOSITY SOLUTIONS TO HAMILTON-JACOBI EQUATIONS ON RIEMANNIAN MANIFOLDS

JUAN FERRERA CUESTA

Departamento de Análisis Matemático. Facultad de Matemáticas. Universidad Complutense de Madrid. España Joint work with D. Azagra and F. López Mesas

Abstract

We establish some perturbed minimization principles, and we develop a theory of subdifferential calculus, for functions defined on Riemannian manifolds. Then we apply these results to show existence and uniqueness of viscosity solutions to Hamilton-Jacobi equations defined on Riemannian manifolds.